OpenStudy (anonymous):

Karen and Jeremy are purchasing a house and are financing $225,000. The mortgage is a 30-year 4/1 ARM at 3.0% with a 2/6 cap structure. What will the remaining balance be after the first 4 years? $196,208.00 $210,667.50 $205,335.00 $210,895.73

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

well, its asking what remaining balance is to refinance after 4 years .. since the 4/1 tells us that it adjusts after 4 years

OpenStudy (amistre64):

my solution strategy is to determine the payments for a fixed 30 year loan, and then balance it out after 4 years of making payments

OpenStudy (amistre64):

what are the payments for the loan?

OpenStudy (anonymous):

I figured it out it is C. $205,335.00

OpenStudy (amistre64):

let me see if i can get that too

OpenStudy (amistre64):

payments i figure are about 948.61 A = Bk^(4*12)-P(1-k^(4*12))/(1-k), for some k compounding factors

OpenStudy (anonymous):

P = B4 = Bk^(4*12)-P(1-k^(4*12))/(1-k) Bk^(4*12)-(Bk^(30*12)(1-k)/(1-k^(30*12)))(1-k^(4*12))/(1-k) 225000k^(4*12)-(225000k^(30*12)(1-k)/(1-k^(30*12)))(1-k^(4*12))/(1-k), k=1+.03/12

OpenStudy (anonymous):

Yes That is how I solved it (:

OpenStudy (amistre64):

yay!! nice formula lol

OpenStudy (anonymous):

THank you so much for your help! :D

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

Well I looked through some of your other responses to other people because I just happened upon them... and you used that formula xp So I was like hmmm well I guess I could try that.

OpenStudy (anonymous):

Because the one that is given to me is difficult... that one made more sense